# Can LLMs Reason with Rules? Logic Scaffolding for Stress-Testing and Improving LLMs

### v2

@@ Abstract
Siyuan Wang1, Zhongyu Wei1, Yejin Choi2,4, Xiang Ren3,4 1Fudan University, 2University of Washington, 3University of Southern California, 4Allen Institute for Artificial Intelligence wangsy18@fudan.edu.cn
@@ 1 Introduction
Humans naturally abstract underlying logic (e.g., inferential rules) from extensive real-world observations Barwise (1993), which is beneficial for addressing diverse reasoning situations. An inferential rule is typically defined as a premise with a set of facts (e.g., “Person X died before … earlier than B”) leading to a conclusion (e.g., ‘‘Person X cannot access Object Y”) Boghossian (2014). Grasping this rule enables the deduction that a person cannot access an object invented posthumously. This work utilizes symbolic logic as a scaffold to generate challenging reasoning situations for GPT-series LLMs. We observe a discernible gap between LLMs and humans in understanding inferential rules, especially rules with complex premises.
@@ 2.2 Primitive Rule Generation Pipeline
Step-1: Conclusion Preparation From the set of abstract objects, we select any two, e.g., Person and Food, and prompt GPT-4 to generate potential predicates connecting them as conclusions, e.g., CanEat(Person X, Food Y). We attempt every possible pairing of two, where the selected objects can be identical. For each pair of objects, {object1} and {object2}, we aim to generate conclusions across five domains: {object affordance, accessibility, interaction, location and person’s need}, thereby covering diverse scenarios. Explanations and example rules of these domains are listed in Appendix A.2. The prompt for conclusion preparation about affordance is below. Besides, we negate the generated predicates to yield both positive and negative conclusions, e.g., CanNotEat(Person X, Food Y), across object affordance, accessibility, and interac-
@@ 3.2 Empirical Analysis
(4) Do LLMs exhibit a polarity bias over inferential rules? Our inferential rules contain both positive and negative conclusions. We conduct a comparative analysis of polarity discrepancy, as shown in Figure 8a. GPT-4 and GPT-3.5-Instruct exhibit a pronounced positive bias, performing better on rules with positive conclusions. This bias may originate from the imbalanced distribution of LLMs’ training data Garg et al. (2022), with a higher proportion of positive statements. We further explore different CoT strategies with GPT-4: (1) first answer then explain (Answer-Explain), (2) first think then answer (Think-Answer), (3) self-consistently think then answer (Self-ConsistencyWang et al. (2022). Various CoT prompts are listed in Appendix B.5. Figure 8b shows that although advanced CoT strategies can mitigate the positive bias, they adversely impact the performance on rules with both positive and negative conclusions.
@@ 4.2 Rule Generation Evaluation
Table 3: LLM evaluation results.
@@ 5 Related Work
Integration of Logical Rules and LLMs Recently, the integration of inferential rules with neural networks, particularly LLMs, has gained significant attention. This approach combines the logical interpretability of symbolic reasoning and adaptive power of neural computing, improving LLMs’ logical reasoning ability.  Wang et al. (2021); Olausson et al. (2023) transform textual statements into logical expressions and conduct symbolic reasoning following logical rules. Weir and Van Durme (2022) train neural models using a set of inferential rules for dynamic application. This direction broadens LLMs’ ability with flexible rule generation and application for complex reasoning.
@@ Acknowledgments
Xiang Ren’s research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200006, the Defense Advanced Research Projects Agency with award HR00112220046, and NSF IIS 2048211.
@@ A.6 Human Verification Templates and Rates
Before human verification, we first craft a qualification task to select AMT annotators from all English-speaking countries (US, UK, New Zealand, Australia, Canada). The prospective workers are presented with three representative test cases and need to predict whether the premise and conclusion are clearly readable, and if the premise logically entails the conclusion. Only those workers correctly passing all the test cases are recruited. The detailed template for human verification is shown as Figure 11. This template is also used for getting human performance in rule probing analysis, wherein a separate cohort of workers is qualified for manual rule probing. Besides, the overall rates of rule acceptance in different domains during human verification are listed Table 9.

### v3

@@ Abstract
Siyuan Wang1, Zhongyu Wei1thanks:   Corresponding author, Yejin Choi2,4, Xiang Ren3,4 1Fudan University, 2University of Washington, 3University of Southern California, 4Allen Institute for Artificial Intelligence wangsy18@fudan.edu.cn
@@ 1 Introduction
Humans naturally abstract underlying logic (e.g., inferential rules) from extensive real-world observations Barwise (1993), which is beneficial for addressing diverse reasoning situations. An inferential rule is typically defined as a premise with a set of facts (e.g., “Person X died before … earlier than B”) leading to a conclusion (e.g., ‘‘Person X cannot access Object Y”) Boghossian (2014). Grasping this rule enables the deduction that a person cannot access an object invented posthumously. This work utilizes symbolic logic as a scaffold to generate challenging reasoning situations for GPT-series LLMs, as shown in Figure 2. We observe a discernible gap between LLMs and humans in understanding inferential rules, especially rules with complex premises.
@@ 2.2 Primitive Rule Generation Pipeline
Step-1: Conclusion Preparation From the set of abstract objects, we select any two, e.g., Person and Food, and prompt GPT-4 to generate potential predicates connecting them as conclusions, e.g., CanEat(Person X, Food Y). We attempt every possible pairing of two, where the selected objects can be identical. For each pair of objects, {object1} and {object2}, we aim to generate conclusions across five domains: {object affordance, accessibility, interaction, location and person’s need}, thereby covering diverse scenarios. Explanations and example rules of these domains are listed in Appendix A.2. The prompt for conclusion preparation about affordance is below. Besides, we negate the generated predicates to yield both positive and negative conclusions, e.g., CanNotEat(Person X, Food Y), across object affordance, accessibility, and interaction domains, building a complete rule set.
@@ 3.2 Empirical Analysis
(4) Do LLMs exhibit a polarity bias over inferential rules? Our inferential rules contain both positive and negative conclusions. We conduct a comparative analysis of polarity discrepancy, as shown in Figure 8a. GPT-4 and GPT-3.5-Instruct exhibit a pronounced positive bias, performing better on rules with positive conclusions. This bias may originate from the imbalanced distribution of LLMs’ training data Garg et al. (2022), with a higher proportion of positive statements. We further explore different CoT strategies with GPT-4: (1) first answer then explain (Answer-Explain), (2) first think then answer (Think-Answer), (3) self-consistently think then answer (Self-ConsistencyWang et al. (2022b). Various CoT prompts are listed in Appendix B.5. Figure 8b shows that although advanced CoT strategies can mitigate the positive bias, they adversely impact the performance on rules with both positive and negative conclusions.
@@ 4.2 Rule Generation Evaluation
Table 3: LLM evaluation results. “Acc”, “Div.”, “Cpx.” and “Abs.” are abbreviations of accuracy, diversity, complexity and abstractness.
@@ 5 Related Work
Integration of Logical Rules and LLMs Recently, the integration of inferential rules with neural networks, particularly LLMs, has gained significant attention. This approach combines the logical interpretability of symbolic reasoning and adaptive power of neural computing, improving LLMs’ logical reasoning ability Wang et al. (2022a); Mu et al. (2023).  Wang et al. (2021) and Olausson et al. (2023) transform textual statements into logical expressions and conduct symbolic reasoning following logical rules. Weir and Van Durme (2022) train neural models using a set of inferential rules for dynamic application. This direction broadens LLMs’ ability with flexible rule generation and application for complex reasoning.
@@ Acknowledgments
Siyuan Wang’s research is supported by National Natural Science Foundation of China (No. 62176058) and National Key R&D Program of China (2023YFF1204800). The project’s computational resources are supported by CFFF platform of Fudan University. Xiang Ren’s research is supported in part by the Office of the Director of National Intelligence (ODNI), Intelligence Advanced Research Projects Activity (IARPA), via the HIATUS Program contract #2022-22072200006, the Defense Advanced Research Projects Agency with award HR00112220046, and NSF IIS 2048211.
@@ A.6 Human Verification Templates and Rates
Before human verification, we first craft a qualification task to select AMT annotators from all English-speaking countries (US, UK, New Zealand, Australia, Canada). The prospective workers are presented with three representative test cases and need to predict whether the premise and conclusion are clearly readable, and if the premise logically entails the conclusion. Only those workers correctly passing all the test cases are recruited. We compensated workers with different rates based on the complexity of the annotated rules, with prices ranging from 0.08 to 0.16 dollars (i.e., 0.08, 0.10, 0.12, 0.14, 0.16) for rules with complexity levels from 1 to 5. On average, this amounts to 0.12 dollars per annotation. Each annotation takes approximately 0.5 minutes to complete, aiming to match a rate of 15 dollars per hour based on their working time.
The detailed template for human verification is shown as Figure 11. This template is also used for getting human performance in rule probing analysis, wherein a separate cohort of workers is qualified for manual rule probing. Besides, the overall rates of rule acceptance in different domains during human verification are listed Table 9.
@@
A.7 Ablation on Rule Filtering
We calculated the rule yield rates for both heuristic filtering and self-critic, which are respectively 77.59% and 80.39%. Together, they totally yield 62.38% of candidate generated rules through the whole rule filtering process. We sample 100 rules before the self-critic and obtain 79 rules after the self-critic, and conduct human verification over them. Before the self-critic, 42 out of 100 rules (42%) are valid. After the self-critic, 40 out of 79 rules (50.63%) are valid. Among the 21 filtered rules, our self-critic can effectively identify 19 invalid rules, thereby enhancing the quality of retained rules.